TGBT14A_912_50 - Various VEGAS tests

Goals

  • Check converter rack connections
  • Check balance problems
  • Back to back VEGAS / KFPA tests
  • X-band RRL tests

Details

IF cables and re-quantization balancing

  • Start at 10:00am. KFPA is not currently at focus
  • Banks A - D have new cables, Banks E - H are currently in an undetermined state (Galen is changing cables).
  • Configure for eos558_vegas, but do not run any scans; run a balance.
  • CM1 is pegged at 0db (cannot remove any more attenuation, but only getting -24dB.
  • CM1 - connected correctly
  • CM5 - connected correctly
  • CM2, CM6 - connected correctly
  • CM3 - was set to 0dB, correct balance would have been 2dB
  • CM7 - connected correctly
  • CM4, CM8 connected correctly
  • Rebalance - get's all of CM1 through CM8 correct, apart from CM1 problem. CM9-16 still disconnected.
  • All of lbw_gain,1 for all eight banks need further adjustment. Investigate these further. Look at Bank B, since A has a CM1 problem.

  • Bank B:
    • adcpwr look good (~ 50 rms) - 20dB
      • filtersnap -26dB
      • filtersnap_nogain -32dB
      • lbw_gain,1 543763383
      • change to 100000000
      • filtersnap now -40dB
  • try again - rebalance - Bank A - D all fail
  • Bank B: through device explorer:
    • measpwr1, -20dB, filtersnap_nogain,pol1,measpwr -32.9dB, filtersnap_pol1,measpwr -26.7dB, lbw_gain,1 556274390 = 87.45 dB
    • need another 6.7dB, = 94.125, = 26000000000 - goes to -12dB
    • 10000000000 - goes to -20dB - doesn't seem correct change, but this is the value we want
    • lbw_gain,1 = 1 nogain = -32dB, gain = -46dB, difference = -14dB = factor of 25 (why not 32)?
    • lbw_gain,1 = 100, nogain = -46dB still!!!!

  • Check Banks E through H connections
  • CM9, CM13 ok, 10, 14 ok, 11,15 ok, 12, 16 ok - all ok

  • ADC balancing again. Balance fails with lbw_gain = 551M, filtersnap = -27.3dB
  • lbw_gain 1000M filtersnap = -22db
  • lbw_gain 2000M filtersnap = -16dB
  • lbw_gain 200M filtersnap = -35dB

Jim Braatz tests

  • Run eso558_vegas and eso558_acs back to back (doing NODs)
  • Scan 7 - 10 VEGAS
  • Scan 11 - 14 ACS
  • Still has the VEGAS spurious last integration problem - Joe will fix - abort these scans
  • Getting an exception error while balancing - Joe will restart turtle
  • Scan 18 - 21 VEGAS - still get a short 6th integration, now has 0 integration time!
  • Joe notes phases are of unequal length, also last integrations, spectrum seems to be copied from previous one. Revert this out.
  • Back to release version
  • Scan 23 26 - VEGAS Same unequal phases - we realize this is due to LO blanking (Doppler tracking) in first phase.
  • Abort and let Joe test some other changes.
  • No useful data obtained

Dana Balser tests

  • Data in AGBT14B_431_05 - Dana is writing a log file which I will ask for.

Topic revision: r6 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback